OpenStudy (anonymous):

Define the variables, write the linear system, the matrix equation, and the answer. There are 27 centipedes and millipedes in a field. Each centipede has 175 pairs of legs and each millipede has 336 pairs of legs. There are 14, 924 legs on the field. How many centipedes are in the field.

OpenStudy (kewlgeek555):

I don't know any of that matrix crap or linear systems, but I knwo the variable must be how many millipede there are OR m millipedes.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):



OpenStudy (wolfe8):

Let x be the number of centipedes and y be the number of millipedes. You are told that each centipede has 175 legs so the number of legs will be 175 times however many centipedes there are, 175x. You are told that there are 27 of them, so 175*27. Then you have a number of millipedes with 336 legs each, making a total of 336y legs. Then you are told that there are 14924 legs in total, so the equation will be (175*27)+336y=14924 Do you know how to write this in matrix form?

OpenStudy (anonymous):

@wolfe8 not exactly, could you help me with that?

OpenStudy (wolfe8):

Tell me what equation you have in the end.

OpenStudy (anonymous):

it's that i dont understand this problem at all..

OpenStudy (wolfe8):

I just explained to you what happened in my first reply...

OpenStudy (anonymous):

ok well im stupid and i need you to dumb this problem really down for me. i know how to matrices but not this one.

OpenStudy (wolfe8):

I don't know how else to say it unless I show you in person... Sorry but this is the best I can do for now

OpenStudy (anonymous):

can't use draw to draw it out?

OpenStudy (anonymous):

you can use the lines and letters on draw to create a matrix. All i really need is for you to create the linear system and the matrix. from there i can finish the problem

OpenStudy (wolfe8):

Alright I will do it one more time since I read the question wrong. S let x=centipede and y=millipede. You are told that there ate 27 centipedes and millipedes, so x+y=27. Then you are told that centipedes have 175 legs each, so total legs for centipedes will be 175x. For millipede you have 336y, correct? And you are told that the total of all those is 14924. So the equation would be 175x+336y=14924 So I showed you how to get the 2 linear equations. You said you could do matrices.

OpenStudy (anonymous):

GRACIAS LOBO♥

OpenStudy (wolfe8):

Hahaha. So you understand it now? And sorry I read the question wrong in my first answer.